# KoopAGRU: A Koopman-based Anomaly Detection in Time-Series using Gated Recurrent Units

### v1

@@ II-B Koopman Theory in Nonlinear Systems Analysis and Prediction
  • Disease Predictions: [58] demonstrated the use of Koopman operator theory in modeling and forecasting the dynamics of epidemiological systems, while [59] expanded this work to further refine disease outbreak predictions using Koopman-based approaches.
  • Sea Ice Prediction: [60] applied Koopman Mode Decomposition (KMD) to satellite data of sea ice concentration, uncovering spatial modes and providing accurate geographic predictions of sea ice concentration up to four years into the future.
@@ A-A The Koopman operator
$[\mathbb{U}\phi](\boldsymbol{z})=\varphi(\mathcal{F}(\boldsymbol{z}))$ (28)
In other terms, we define the Koopman operator, $\mathbb{U}$, as an infinite-dimensional linear operator that pairs up with the dynamical system and acts on any scalar-valued observable function $\phi$ within $\mathcal{F}(M)$ that belongs to the system. The action of $\mathbb{U}$ on $\phi$ is given by $[\mathbb{U}\phi](z)=\varphi(\mathcal{F}(z))$ and the evolution of these observable functions is facilitated by $\mathbb{U}$. A noteworthy property of the operator $\mathbb{U}$ is that it maintains the linearity and invariance of the function space $\mathcal{F}(M)$.
@@ A-B DeepDMD
In deepDMD, the function $\psi$ plays a crucial role as it transfers time-series data into an observable space. However, the task does not end there, as the approach also requires an inverse mapping that enables the retrieval of the original measurements from the observable space. Two approaches can accomplish this reversal. The first method involves identifying a function $\phi$, which is essentially the reverse of $\psi$. When this $\phi$ function is recognized, it suggests that $X_{t}$ can equivalently be depicted as $\phi(\psi(X_{t}))$. In contrast, the reverse operation can be carried out by conjoining the actual measured states, represented by $X_{t}$, with observables to achieve measurement-inclusive observables. This amalgamation results in a composite function $\Psi(X_{t})$ that incorporates both $X_{t}$ and $\psi(X_{t})$.

### v2

@@ II-B Koopman Theory in Nonlinear Systems Analysis and Prediction
  • Disease Predictions [58] demonstrated the use of Koopman operator theory in modeling and forecasting the dynamics of epidemiological systems, while [59] expanded this work to further refine disease outbreak predictions using Koopman-based approaches.
  • Sea Ice Prediction [60] applied Koopman Mode Decomposition (KMD) to satellite data of sea ice concentration, uncovering spatial modes and providing accurate geographic predictions of sea ice concentration up to four years into the future.
@@ A-A The Koopman operator
$[\mathbb{U}\varphi](\boldsymbol{z})=\varphi(\mathcal{F}(\boldsymbol{z}))$ (28)
In other terms, we define the Koopman operator, $\mathbb{U}$, as an infinite-dimensional linear operator that pairs up with the dynamical system and acts on any scalar-valued observable function $\varphi$ within $\mathcal{F}(M)$ that belongs to the system. The action of $\mathbb{U}$ on $\varphi$ is given by $[\mathbb{U}\varphi](z)=\varphi(\mathcal{F}(z))$ and the evolution of these observable functions is facilitated by $\mathbb{U}$. A noteworthy property of the operator $\mathbb{U}$ is that it maintains the linearity and invariance of the function space $\mathcal{F}(M)$.
@@ A-B DeepDMD
In deepDMD, the function $\psi$ plays a crucial role as it transfers time-series data into an observable space. However, the task does not end there, as the approach also requires an inverse mapping that enables the retrieval of the original measurements from the observable space. Two approaches can accomplish this reversal. The first method involves identifying a function $\varphi$, which is essentially the reverse of $\psi$. When this $\varphi$ function is recognized, it suggests that $X_{t}$ can equivalently be depicted as $\varphi(\psi(X_{t}))$. In contrast, the reverse operation can be carried out by conjoining the actual measured states, represented by $X_{t}$, with observables to achieve measurement-inclusive observables. This amalgamation results in a composite function $\Psi(X_{t})$ that incorporates both $X_{t}$ and $\psi(X_{t})$.